At what annual rate would the following have to be invested?
$5,143, to grow to $44,972, in 40 years.

Round the answer to two decimal places in percentage form
Answer #1

Future value = Present value ( 1+r)^n

44972 = 5143 (1+R)^N

44972 = 5143 ( 1+R)^40

44972 /5143 = (1+R) ^40

8.744 = (1+R)^40

R = 5.57%